OpenStudy (anonymous):

Show that the parametric equations x=-e^2t and y=e^t are equal to the Cartesian equation x^2-y^4= 0when -inf

OpenStudy (anonymous):

x^2-y^4 =0 ?

OpenStudy (anonymous):

Yes I forgot to put that in

OpenStudy (anonymous):

all right , sub x=-e^2t and y=e^t into the equation

OpenStudy (anonymous):

How could you plug them into x^2-y^4 when you are trying show that?

OpenStudy (perl):

you are showing that (or proving that) IF x = -e^(2t) and y = e^t THEN x^2 -y^4 = 0

OpenStudy (perl):

and the latter is a cartesian equation

OpenStudy (anonymous):

x=- e^2t x^2 =-(e^2t)^2= - e^4t

OpenStudy (perl):

x^2 = e^(4t) ,

OpenStudy (anonymous):

Where did the negative sign go when you squared -e^(2t)

OpenStudy (perl):

when you square a negative it becomes positive

OpenStudy (anonymous):

So you can just include the negative as part of the number.

OpenStudy (perl):

im not sure what youre asking

OpenStudy (perl):

Claim: IF x = -e^(2t) and y = e^t THEN x^2 -y^4 = 0 Proof: Suppose x = -e^(2t) and y = e^t then by substitution x^2 - y^4 = (-e^(2t))^2 - (e^t)^4 = e^(4t) - e^(4t) = 0

OpenStudy (anonymous):

I see what you are saying

OpenStudy (perl):

but it is not obvious how they got this cartesian equation